Find out the ratio of velocities of the given signal in the core and cladding if the refractive indices are 1.5 and 1.48 respectively

1
Expert's answer
2022-01-02T18:23:12-0500

"v=\frac cn,"

"\frac{v_2}{v_1}=\frac{n_2}{n_1}=\frac{1.50}{1.48}=1.014."